Wednesday 03/26/14

9:20a – 10:10p,

remotely, 10:50p – 11:30p


  • passage cells
  • washout probe from new multicolor stain
  • finish updating notes from yesterday
    • STORM run configurations
    • data analysis launched
  • break down PRISM2 run of L3C04. (complete)
  • break down STORM2 run of L3C03 (complete)
  • set up copying and data-fitting of L3C03 and L3C04 data (copied, started fitting C03)
  • start making new probes from Lib2: F01, F02, F07, G06, G08, G09

Longer goals

  • More probe making
    • lower priority, remakes: F06, G05,
    • Lib3 on deck: embedded yellow regions E03-E10
  • start ChromatinCropper of L3C01
  • ChromatinCropper of F05-F06 02/28 data (and 02/27 data?)
  • analyze buffer data



New multicolor stains

  • New stains failed. I believe I previously verified the low concentration (1.5 uL probe) of these probes makes this happen. Repeat at higher concentration.

Sample C05 on STORM2

  • 3 uL COT, 5 uL BME buffer
  • staining looks pretty good, spots look decent.

Sample C06 on STORM4

  • 5 uL COT, 5 uL BME.
  • dots are clearly fainter
  • switching pretty well, background is tolerable but a bit higher than desired for imaging small (<50kb) chromatin regions.

Cell Staining

  • Need to stain new cells today if I want to take adventage of the microscope time I have this week.
  • Despite just passaging cells, between the two culture flasks I think I easily have enough cells to plate another 12 coverslips. Let’s fix new cells tomorrow.
  • New stains: C07, C08, and double stain F06-A647 + F05-A750.

New probes making

  • Lib2: F01, F02, F03, F05, F06, F07, G06, G08
  • ran PCR. F05 sublibrary failed, rest look good. Press ahead with all
  • ran PCR clean up / DNA clean and concentrate 5.
  • setup 20 uL T7 reactions to run overnight.


Data Analysis

  • all data running last night finished except the 02/28 double color data (which I was excited to analyze). This is still chugging along at a conservative 25% CPU on Cajal.
  • canceled and relaunched analysis with 8 parallel processes per color. Along with the C03 analysis this has Cajal up to 100% CPU engagement overnight again.

Ph project

  • email Nicole clarification and explanation of graphs.
  • further email discussion and clarifications
  • To do:
    1. normalize all distribution plots. Report both total Ph localizations and total clusters.
    2. consider truncating distributions at 700 rather than clustering into last bin. It’s important this doesn’t look like a second peak.
    3. add full length distributions (out to 1.5 um) into supplement.
This entry was posted in Summaries and tagged , , , , . Bookmark the permalink.